# Computing Certificates of Strictly Positive Polynomials in Archimedean Quadratic Modules

### v1

@@ 5 Computing certificates using extended Averkov construction and Lasserre’s method
Let $m=\deg(f)$ and consider the sum of squares
$\delta=c\|\bm{x}\|^{2\lfloor\frac{m}{2}\rfloor+2},\mbox{ where }c=\frac{1}{2}\min_{x\in\operatorname{S}(g)}\frac{f(\bm{x})}{\|\bm{x}\|^{2\lfloor\frac{m}{2}\rfloor+2}g(\bm{x})}.$ (5)
Next we prove that $\bar{f}:=f-\delta g$ is strictly positive on $\operatorname{S}(g)$ and has a lower bound over $\mathbb{R}^{n}$. Suppose that $f$ is in the form $f=\sum_{\bm{\alpha}}\lambda_{\bm{\alpha}}\bm{x}^{\bm{\alpha}}$ with $\lambda_{\bm{\alpha}}\in\mathbb{R}$. Then for any $\bar{\bm{x}}\in\operatorname{S}(g)$, we have $\frac{f(\bar{\bm{x}})}{\|\bm{x}\|^{2\lfloor\frac{m}{2}\rfloor+2}g(\bar{\bm{x}})}>c$ and thus we have $\bar{f}(\bar{\bm{x}})=f(\bar{\bm{x}})-\delta(\bar{\bm{x}})g(\bar{\bm{x}})>0$.
In the first paragraph of the proof of Theorem 3.3, we show that there exists $K>0$ such that $\operatorname{S}(g+K)$ is bounded. Then for any $\bm{\bar{x}}\notin\operatorname{S}(g+K)$, we have $g(\bm{x})<-K<0$, and thus $|f|\leq\sum_{\bm{\alpha}}|\lambda_{\bm{\alpha}}\bm{x}^{\alpha}|\leq\sum_{\bm{\alpha}}|\lambda_{\bm{\alpha}}|\cdot\|\bm{x}\|^{|\bm{\alpha}|}$. Now consider the rational function
$\left|\frac{f}{\delta g}\right|\leq\left|\frac{f}{c\|\bm{x}\|^{2\lfloor\frac{m}{2}\rfloor+2}g}\right|\leq\sum_{\bm{\alpha}}\frac{|\lambda_{\bm{\alpha}}|}{cK}\|\bm{x}\|^{|\bm{\alpha}|-2\lfloor\frac{m}{2}\rfloor-2}.$
Note that $m=\deg(f)$, therefore $|\bm{\alpha}|-2\lfloor\frac{m}{2}\rfloor-2<0$ for any $\bm{\alpha}$. When $\|\bm{x}\|\to\infty$, we have $\|\bm{x}\|^{|\bm{\alpha}|-2\lfloor\frac{m}{2}\rfloor}\to 0$, and thus there exists $M>0$ such that $|\frac{f}{\delta g}|\leq 1$ for any $\bm{x}$ such that $\|\bm{x}\|\geq M$. In other words, for any $\bar{\bm{x}}\in\overline{\operatorname{S}(g+K)}\cap\{\bar{\bm{x}}\leavevmode\nobreak\ |\leavevmode\nobreak\ \|\bm{x}\|\geq M\}$, we have $\frac{f}{\delta g}(\bar{\bm{x}})-1<0$. Note that $\bar{\bm{x}}\not\in\operatorname{S}(g)$, and thus $g(\bar{\bm{x}})<0$. Then we have $\bar{f}(\bar{\bm{x}})=f(\bar{\bm{x}})-\delta(\bar{\bm{x}})g(\bar{\bm{x}})>0$. This indicates that the set $\operatorname{S}(-\bar{f})$ is contained in $\{\|\bm{x}\|\geq M\}$ and bounded, and thus $\bar{f}$ has a lower bound over $\mathbb{R}^{n}$. ∎
@@ 5.2 Algorithm for solving the certificate problem
Integrating the extended Averkov construction and Lasserre’s method, the result is the following steps which is later formulated as Algorithm 4.
@@
    If $\tilde{f}$ does not have a lower bound over $\mathbb{R}^{n}$ (otherwise we can skip the construction here and set $\hat{f}=\tilde{f}$), construct a sum of squares $\delta$ in (5) such that, as shown by Proposition 5.1, $\hat{f}:=\tilde{f}-\delta g\in\operatorname{QM}(g)$ is strictly positive on $\operatorname{S}(g)$ and has a lower bound over $\mathbb{R}^{n}$.
@@
    Applying Theorem 3.5 (Algorithm 2) to $\hat{f}$ for $B=\mathbb{R}^{n}$, construct a polynomial $\hat{g}\in\operatorname{QM}(G)$ such that $\hat{f}-\hat{g}>0$ over $\mathbb{R}^{n}$.
@@
    In the univariate case when $n=1$, $\hat{f}-\hat{g}$ is already a sum of squares itself.
    $f=\underbrace{\hat{f}-\hat{g}}_{\text{sum of squares}}+\underbrace{\hat{g}+\delta g+\tilde{g}}_{\in\operatorname{QM}(g)}\in\operatorname{QM}(G).$
@@
    $f=\underbrace{\hat{f}-\epsilon p_{r}\hat{g}}_{\text{sum of squares}}+\underbrace{\epsilon p_{r}\hat{g}+\delta g+\tilde{g}}_{\in\operatorname{QM}(g)}\in\operatorname{QM}(G),$
    and the cerficates of $f$ w.r.t. $\operatorname{QM}(G)$ also follow from combination of those of $\hat{g}$, $g$, and $\tilde{g}$ w.r.t. $\operatorname{QM}(G)$
Input: A polynomial set $G=\{g_{1},\dots,g_{s}\}\subseteq\mathbb{R}[x]$ and a polynomial $f\in\mathbb{R}[x]$ such that $f>0$ over $\operatorname{S}(G)$
Output: A sequence $(\sigma_{0},\sigma_{1},\dots,\sigma_{s+1})$ of sum-of-squares multipliers such that $f=\sigma_{0}+\sum_{i=1}^{s}{\sigma_{i}g_{i}}+\sigma_{s+1}(N-\sum_{i=1}^{s}x_{i}^{2})$
@@ 14             Construct $c_{i}$ and $c_{j}$ as in (4);
18$(\sigma_{1},\ldots,\sigma_{s}):=\operatorname{\sf Averkov}(G,f,\operatorname{S}(g))$;
@@ 19 $\tilde{f}:=f-\sum_{i=1}^{s}\sigma_{i}g_{i}$;
21       Construct $\delta$ as in (5); $\tilde{f}:=\tilde{f}-\delta g$;
22
@@ 26       $\sigma_{i}:=\sigma_{i}+(p+\delta)c_{i}$; $\sigma_{j}:=\sigma_{j}+(p+\delta)c_{j}$;
27       $\sigma_{0}:=f-\sum_{i=1}^{s}{\sigma_{i}g_{i}}$;
28
29else
30       $q:=\operatorname{\sf ExtLasserre}(\tilde{f}-pg,g)$ ;
@@ 32             $\sigma_{s+1}=(q+p+\delta)c_{s+1}$ ;
35             $\sigma_{i}=\sigma_{i}+(p+q+\delta)c_{i}$; $\sigma_{j}=\sigma_{j}+(p+q+\delta)c_{j}$;
36
37      $\sigma_{0}:=f-\sum_{i=1}^{s+1}{\sigma_{i}g_{i}}$;
38
@@ 37      $\sigma_{0}:=f-\sum_{i=1}^{s+1}{\sigma_{i}g_{i}}$;
40return $(\sigma_{0},\sigma_{1},\dots,\sigma_{s+1})$;
41
@@
Now let us illustrate Algorithm 4 with the univariate quadra-tic module $\operatorname{QM}(g_{1},g_{2})$, where
@@
Then we have $\operatorname{S}(g_{1},g_{2})=\{0,1,2\}$. The polynomial
@@
$g:=g_{1}+x^{2}g_{2}=-\frac{1}{2}x(x-1)(x-2)(3x-1)$
in $\operatorname{QM}(g_{1},g_{2})$ using (4), and now we have $\operatorname{S}(g)=[0,\frac{1}{3}]\cup[1,2]$. Since $f>0$ on $\operatorname{S}(g)$, we can skip the Averkov’s method (Lemma 2.3). Now, construct the polynomial $\delta=20x^{4}$ by using (5) in line 14 of Algorithm 4. The polynomial $\tilde{f}=f-\delta g$ has a global lower bound $-159$. In line 15, by using Algorithm 2, we have $\bar{f}:=\tilde{f}-\sigma g>0$ on $\mathbb{R}$ so it is a sum of squares in $\mathbb{R}[x]$, where
@@
Consequently, Algorithm 4 terminates with the certificate $f=(\delta+\sigma)g_{1}+(\delta+\sigma)x^{2}g_{2}+\tilde{f}$.
To provide a comparison with the alternative method that does not construct the polynomial $\tilde{f}$ in lines 13-14 but uses Theorem 3.3 directly, we present the certificates computed by this alternative method. We have $\bar{f}:=f-\bar{\sigma}g>0$ on $\mathbb{R}$ so it is a sum of squares in $\mathbb{R}[x]$, where
@@
Finally, writing $g$ in terms of $\{g_{1},g_{2}\}$, we have the certificates of $f$ w.r.t. $\{g_{1},g_{2}\}$: $f=\sigma g_{1}+\sigma x^{2}g_{2}+\bar{f}$ and the certificates computed in this way have higher degrees than those by Algorithm 4.
@@ 6 Experiments
This section reports the experimental results with our prototypical implementations  [^2] The implementation can be found using the following link: https://github.com/typesAreSpaces/StrictlyPositiveCert in the Computer Algebra System Maple to compute certificates of strictly positive polynomial over bounded semialgebraic sets on a variety of problems. We implement the extended Lassere method discussed in Section 4 using the Julia programming language  [^3] The implementation is hosted in GitHub and can be accessed using the following link: https://github.com/typesAreSpaces/globalStrictPositive. We also compare our results with RealCertify [8], a Maple package that computes certificates of strictly positive elements in quadratic modules using a semidefinite programming approach. All experiments were performed on an M1 MacBook Air with 8GB of RAM running macOS.

### v2

@@ 6 Computing certificates using the extended Averkov construction and Lasserre’s method
Let $m=\deg(f)+(\deg(g)-1)^{n}-\deg(g)$ and consider the sum of squares
$\delta=c\|\bar{\bm{x}}\|^{2\lfloor\frac{m}{2}\rfloor+2},\mbox{ where }c<\frac{1}{2}\min_{\bar{\bm{x}}\in\operatorname{S}(g)}\frac{f(\bar{\bm{x}})}{\|\bar{\bm{x}}\|^{2\lfloor\frac{m}{2}\rfloor+2}g(\bar{\bm{x}})}.$ (9)
Next, we prove that $\bar{f}:=f-\delta\cdot g$ is strictly positive on $\operatorname{S}(g)$ and has a lower bound over $\mathbb{R}^{n}$. Suppose that $f$ is of the form $f=\sum_{\bm{\alpha}}\lambda_{\bm{\alpha}}\bm{x}^{\bm{\alpha}}$ with $\lambda_{\bm{\alpha}}\in\mathbb{R}$. Then for any $\bar{\bm{x}}\in\operatorname{S}(g)$, we have $\frac{f(\bar{\bm{x}})}{\|\bar{\bm{x}}\|^{2\lfloor\frac{m}{2}\rfloor+2}g(\bar{\bm{x}})}>c$ and thus we have $\bar{f}(\bar{\bm{x}})=f(\bar{\bm{x}})-\delta(\bar{\bm{x}})g(\bar{\bm{x}})>0$.
Analogously, in the first two paragraphs of the proof of ?THM? LABEL:lemma:d, there exist $l>0$ and $R\geq 1$ such that $|\|\bar{\bm{x}}\|^{(\deg(g)-1)^{n}}g(\bar{\bm{x}})|\geq l\|\bar{\bm{x}}\|^{\deg(g)}$ for $\|\bar{\bm{x}}\|>R$. Then for any $\|\bar{\bm{x}}\|>R$, we have
$\begin{split}|\|\bar{\bm{x}}\|^{2\lfloor\frac{m}{2}\rfloor+2}g(\bar{\bm{x}})|\geq|\|\bar{\bm{x}}||^{m+1}g(\bar{\bm{x}})|\geq l\|\bar{\bm{x}}\|^{\deg(g)-(\deg(g)-1)^{n}+m+1}=l\|\bar{\bm{x}}\|^{\deg(f)+1},\end{split}$
and $|f|\leq\sum_{\bm{\alpha}}|\lambda_{\ p{\alpha}}\bm{x}^{\alpha}|\leq\sum_{\bm{\alpha}}|\lambda_{\bm{\alpha}}|\cdot\|\bar{\bm{x}}\|^{|\bm{\alpha}|}$.
Now consider the rational function
$\left|\frac{f}{\delta\cdot g}\right|\leq\left|\frac{f}{c\|\bar{\bm{x}}\|^{2\lfloor\frac{m}{2}\rfloor+2}\cdot g}\right|\leq\sum_{\bm{\alpha}}\frac{|\lambda_{\bm{\alpha}}|}{c\cdot l}\|\bar{\bm{x}}\|^{|\bm{\alpha}|-\deg(f)-1}.$
Note that $|\bm{\alpha}|-\deg(f)-1<0$ for any $\bm{\alpha}$. When $\|\bar{\bm{x}}\|\to\infty$, we have $\|\bar{\bm{x}}\|^{|\bm{\alpha}|-\deg(f)-1}\to 0$, and thus there exists $M>0$ such that $|\frac{f}{\delta\cdot g}|\leq 1$ for any $\bar{\bm{x}}$ such that $\|\bar{\bm{x}}\|\geq M$. In other words, for any $\bar{\bm{x}}\in\overline{\operatorname{S}(g)}\cap\{\bar{\bm{x}}\penalty 10000\ |\penalty 10000\ \|\bar{\bm{x}}\|\geq M\}$, we have $\frac{f}{\delta\cdot g}(\bar{\bm{x}})-1<0$. Note that $\bar{\bm{x}}\not\in\operatorname{S}(g)$, and thus $g(\bar{\bm{x}})<0$. Then we have $\bar{f}(\bar{\bm{x}})=f(\bar{\bm{x}})-\delta(\bar{\bm{x}})g(\bar{\bm{x}})>0$. This indicates that the set $\operatorname{S}(-\bar{f})$ is contained in $\{\|\bar{\bm{x}}\|\geq M\}$ and bounded, and thus $\bar{f}$ has a lower bound over $\mathbb{R}^{n}$. ∎
@@ 6.2 Algorithm for solving the certificate problem
Integrating the extended Averkov construction and Lasserre’s method, the result is the following algorithm for computing the certificates in Archime-dean quadratic modules, which are later formulated as Algorithm 4.
@@
    If $\tilde{f}$ does not have a lower bound over $\mathbb{R}^{n}$ (otherwise we can skip the construction here and set $\hat{f}=\tilde{f}$), construct a sum of squares $\delta$ as in (9) such that, as shown by ?THM? LABEL:prop:lowerBound, $\hat{f}:=\tilde{f}-\delta\cdot g\in\operatorname{QM}(g)$ is strictly positive on $\operatorname{S}(g)$ and has a lower bound over $\mathbb{R}^{n}$.
@@
    Applying ?THM? LABEL:cor:lowerbound (Algorithm 2) to $\hat{f}$ for $B=\mathbb{R}^{n}$, construct a polynomial $\hat{g}\in\operatorname{QM}(G)$ such that $\hat{f}-\hat{g}>0$ over $\mathbb{R}^{n}$.
@@
    In the univariate case when $n=1$, $\hat{f}-\hat{g}$ is already a sum of squares itself and we have
    $f=\underbrace{\hat{f}-\hat{g}}_{\text{sum of squares}}+\underbrace{\hat{g}+\delta\cdot g+\tilde{g}}_{\in\operatorname{QM}(g)}\in\operatorname{QM}(G).$
@@
    $f=\underbrace{\hat{f}-\epsilon p_{r}\cdot\hat{g}}_{\text{sum of squares}}+\underbrace{\epsilon p_{r}\cdot\hat{g}+\delta\cdot g+\tilde{g}}_{\in\operatorname{QM}(g)}\in\operatorname{QM}(G),$
    and the cerficates of $f$ w.r.t. $\operatorname{QM}(G)$ also follow from combination of those of $\hat{g}$, $g$, and $\tilde{g}$ w.r.t. $\operatorname{QM}(G)$.
Input: A polynomial set $G=\{g_{1},\dots,g_{s}\}\subseteq\mathbb{K}[X]$ and a polynomial $f\in\mathbb{K}[X]$ such that $f>0$ over $\operatorname{S}(G)$
[…]
@@ 15    $g:=c_{i}\cdot g_{i}+c_{j}\cdot g_{j}$;
18
[…]
@@ 20 $\tilde{f}:=f-\sum_{i=1}^{s}\sigma_{i}g_{i}$;
22   Construct $\delta$ as in (9); $\tilde{f}:=\tilde{f}-\delta\cdot g$;
23
@@ 27 $\sigma_{i}:=\sigma_{i}+(p+\delta)c_{i}$; $\sigma_{j}:=\sigma_{j}+(p+\delta)c_{j}$;
28 $\sigma_{0}:=f-\sum_{i=1}^{s}{\sigma_{i}\cdot g_{i}}$;
29
30else
31 $q:=\operatorname{\sf ExtLasserre}(\tilde{f}-p\cdot g,g)$ ;
@@ 33    $\sigma_{s+1}=(q+p+\delta)c_{s+1}$ ;
36    $\sigma_{i}=\sigma_{i}+(p+q+\delta)c_{i}$;
37
38 $\sigma_{0}:=f-\sum_{i=1}^{s+1}{\sigma_{i}\cdot g_{i}}$;
@@
40
41return $(\sigma_{0},\sigma_{1},\dots,\sigma_{s+1})$;
42
@@
Let us illustrate Algorithm 4 with the univariate quadratic module $\operatorname{QM}(g_{1},g_{2})$, where
@@
Then, we have $\operatorname{S}(g_{1},g_{2})=\{0,1,2\}$. The polynomial
@@
$g:=g_{1}+x^{2}\cdot g_{2}=-\frac{1}{2}x(x-1)(x-2)(3x-1)$
in $\operatorname{QM}(g_{1},g_{2})$ using (8), and now we have $\operatorname{S}(g)=[0,\frac{1}{3}]\cup[1,2]$. Since $f>0$ on $\operatorname{S}(g)$, we can skip the Averkov’s method (?THM? LABEL:lemma:c). Now, construct the polynomial $\delta=20x^{4}$ by using (9) in line 4 of Algorithm 4. The polynomial $\tilde{f}=f-\delta\cdot g$ has a global lower bound 159-159. In line 4, by using Algorithm 2, we have $\bar{f}:=\tilde{f}-\sigma\cdot g>0$ on $\mathbb{R}$ so it is a sum of squares in $\mathbb{Q}[x]$, where
@@
Consequently, Algorithm 4 terminates with the certificate $f=(\delta+\sigma)\cdot g_{1}+(\delta+\sigma)x^{2}\cdot g_{2}+\tilde{f}$.
To provide a comparison with the alternative method that does not construct the polynomial $\tilde{f}$ with a global lower bound in lines 44 but uses ?THM? LABEL:lemma:d directly, we present the certificates computed by this alternative method. We have $\bar{f}:=f-\bar{\sigma}\cdot g>0$ on $\mathbb{R}$ so it is a sum of squares in $\mathbb{K}[X]$, where
@@
Finally, writing $g$ in terms of $\{g_{1},g_{2}\}$, we have the certificates of $f$ w.r.t. $\{g_{1},g_{2}\}$: $f=\sigma\cdot g_{1}+\sigma x^{2}\cdot g_{2}+\bar{f}$ and the certificates computed in this way have higher degrees than those by Algorithm 4.
Example 6.3.
Let us consider the another example in Schmüdgen (2025, Lemma 18). Let the quadratic module $\operatorname{QM}(G)=\operatorname{QM}(g_{1},g_{2},g_{3},g_{4})\subseteq\mathbb{Q}[x,y]$, where $g_{1}=x,g_{2}=y,g_{3}=(1-x)(1-y)\text{ and }g_{4}=2-(x+y).$ We have $\operatorname{S}(G)=[0,1]^{2}$. Let
f=(x+1)(2x)+(y+1)(2y),f=(x+1)(2-x)+(y+1)(2-y),
which is strictly positive on $\operatorname{S}(G)$, and thus $f\in\operatorname{QM}(G)$ by Putinar’s Positivstellensatz.
Now consider the first step to construct the polynomial $g$. Since $2-x=g_{4}+g_{1}\in\operatorname{QM}(G)$, we have $(2-x)x=\frac{x^{2}}{2}(2-x)+2(1-\frac{x}{2})^{2}x\in\operatorname{QM}(G)$. Then $4-x^{2}=2(2-x)+(2-x)x\in\operatorname{QM}(G)$. Similarly, we have $4-y^{2}\in\operatorname{QM}(G)$. To sum up, we assign $g=8-x^{2}-y^{2}=(2+\frac{x^{2}}{2}+2(1-\frac{x}{2})^{2})\cdot g_{1}+(2+\frac{y^{2}}{2}+2(1-\frac{y}{2})^{2})\cdot g_{2}+(4+\frac{x^{2}}{2}+\frac{y^{2}}{2})\cdot g_{4}\in\operatorname{QM}(G)$, which is an Archimedean polynomial.
Then, for the subset $B=\operatorname{S}(g)=\{\|\bar{\bm{x}}\|\leq 2\sqrt{2}\}$ we construct sums of squares $\sigma_{i}\penalty 10000\ (i=1,\ldots,4)$ by using Averkov’s method, such that $\tilde{f}=f-\sum_{i=1}^{4}\sigma_{i}\cdot g_{i}>0$ over $B$, where
$\sigma_{i}=\left(\frac{10}{51}\cdot g_{i}-\frac{50}{51}\right)^{4}.$
Since $\tilde{f}$ has no lower bound, we construct the polynomial $\delta=\frac{(x^{2}+y^{2})^{5}}{6000}$ by using (9) in line 4 of Algorithm 4. The polynomial $\tilde{f}=\tilde{f}-\delta\cdot g$ is already sums of squares, hence the algorithm terminates. Finally, we have the certificate of $f$ in $\operatorname{QM}(G)$: $f=\sum_{i=1}^{4}\sigma_{i}\cdot g_{i}+\delta\cdot g+\tilde{f}$.
In general, the polynomial $\tilde{f}$ at line 4 is not sums of squares. For example, when construct $\delta=\frac{(x^{2}+y^{2})^{5}}{40000}$, the polynomial $\tilde{f}$ has a global lower bound of 6-6. In line 4, by using Algorithm 2, we have $\bar{f}:=\tilde{f}-\sigma\cdot g>0$ on $\mathbb{R}^{2}$, and is also a sums of squares, where
$\sigma=\left(\frac{10}{47}\cdot g-\frac{40}{47}\right)^{6}.$
Consequently, Algorithm 4 terminates with the certificate $f=\sum_{i=1}^{4}\sigma_{i}\cdot g_{i}+(\delta+\sigma)\cdot g+\bar{f}$.
@@ 7 Experiments
This section reports the experimental results with our prototypical implementations  [^3] The implementation can be found using the following link: https://github.com/typesAreSpaces/StrictlyPositiveCert in the Computer Algebra System Maple to compute certificates of strictly positive polynomial over bounded semialgebraic sets on a variety of problems. We implemented the extended Lassere’s method discussed in Section 4 using the Julia programming language  [^4] The implementation is hosted in GitHub and can be accessed using the following link: https://github.com/typesAreSpaces/globalStrictPositive. We also compare our results with RealCertify Magron and Safey El Din (2018b), a Maple package that computes certificates of strictly positive elements in quadratic modules using a semidefinite programming approach. All experiments were performed on an M1 MacBook Air with 8GB of RAM running macOS.